OpenStudy (anonymous):

Please Help! Use the Pythagorean theorem to find one leg if the hypotenuse is 13 and the other leg is 5. 12 3 13.93 11.24

OpenStudy (anonymous):

13^2+5^2=c^2 169+25=c^2 194=c^2 √194=c 13.9283883=c

OpenStudy (anonymous):

\[A ^{2}+B ^{2}=C ^{2}\] Hypotenuse = C A and B are your other legs. So plug 13 in for C, 5 in for either A or B and solve for the remaining variable.

OpenStudy (anonymous):

Thanks:)